# JointDiT: Enhancing RGB-Depth Joint Modeling with Diffusion Transformers

### v2

@@ Abstract
We present JointDiT, a diffusion transformer that models the joint distribution of RGB and depth. By leveraging the architectural benefit and outstanding image prior of the state-of-the-art diffusion transformer, JointDiT not only generates high-fidelity images but also produces geometrically plausible and accurate depth maps. This solid joint distribution modeling is achieved through two simple yet effective techniques that we propose, i.e., adaptive scheduling weights, which depend on the noise levels of each modality, and the unbalanced timestep sampling strategy. With these techniques, we train our model across all noise levels for each modality, enabling JointDiT to naturally handle various combinatorial generation tasks, including joint generation, depth estimation, and depth-conditioned image generation by simply controlling the timestep of each branch. JointDiT demonstrates outstanding joint generation performance. Furthermore, it achieves comparable results in depth estimation and depth-conditioned image generation, suggesting that joint distribution modeling can serve as a replaceable alternative to conditional generation. The project page is available at https://byungki-k.github.io/JointDiT/
Figure 1: We present JointDiT that models the joint distribution of RGB images and depth maps. We visualize the jointly generated images, depth maps, and their 3D lifting results of JointDiT. By leveraging the strong image prior of the state-of-the-art diffusion transformer [5], JointDiT generates high-fidelity images and geometrically plausible and accurate depth maps.
$\dagger$\$$\dagger$\$footnotetext: Work done during an internship at Microsoft Research Asia.
Figure 2: 3D lifting results of our JointDiT and depth diffusion models, i.e., Marigold [28] and Geowizard [18]. JointDiT shows plausible and smooth 3D point clouds than depth estimation models because our joint distribution model inherently learns the relationship between images and depth across repetitive generative processes. In contrast, depth estimation models exhibit rough surfaces due to the uncertainty in depth estimation.
@@ 1 Introduction
In the era of generative AI, diffusion models have made remarkable advancements in synthesizing images [23, 45]. The outstanding capability of text-to-image diffusion models has been found to be useful not only for image generation but also for solving important inverse problems [10, 11, 9], image inpainting [12, 37], image editing [27, 13], and even further cross-modal conditional generation, such as depth-conditioned image generation [61, 4, 38], and image-conditioned depth estimation [28, 18, 19]. These works have demonstrated that utilizing the image prior of diffusion models is effective for modeling conditional distribution.
Recently, in image and depth modalities, joint distribution modeling [53, 60, 31] has shown that joint distribution modeling not only enables joint generation but also shows a potential as a replaceable alternative to existing depth estimation methods and depth-conditional image generation methods within a single unified framework by treating them as special cases of joint modeling. It demonstrates that joint modeling can be easily generalized for various tasks including controllable and conditional generation and estimation. While the joint modeling approaches have shown their rich versatility, the realism of the generation is limited.
@@
To achieve this, we model the joint distribution by harnessing the strong image prior of state-of-the-art diffusion transformer [5] and building a parallel depth branch through joint connection modules. By training on separate noise levels for each modality, JointDiT flexibly facilitates combinatorial tasks of image and depth, including the joint task, by simply controlling the timestep of each branch. To enable separate noise level training, we also propose two simple yet effective techniques, i.e., adaptive scheduling weights and unbalanced timestep sampling strategy, designed for multi-modal diffusion training with separate noise levels. As shown in Figure 2, JointDiT produces smoother, more structured, and more plausible 3D lifting results compared to conditional generation models [18, 28], which exhibit rough surfaces due to the uncertainty in depth estimation. JointDiT also achieves significantly superior joint generation results compared to previous joint generation methods [53, 60, 31] while demonstrating comparable performance in conditional generation tasks, such as depth estimation and depth-conditioned image generation.
@@
  • We present JointDiT, a model for solid joint distribution modeling between image and depth modalities across all noise levels by leveraging the strong image prior of diffusion transformers. It flexibly facilitates combinatorial tasks, such as joint generation, depth estimation, and depth-conditioned image generation by simply controlling the timestep of each branch.
  • We propose adaptive scheduling weights and the unbalanced timestep sampling strategy for separate noise level training in multi-modality, which significantly contribute to improving the performance of combinatorial tasks. Through these techniques, we demonstrate that joint distribution modeling can be the replaceable alternative of conditional generation.
Figure 3: Overall pipeline of JointDiT. Building on Flux [5], we introduce a parallel depth branch with trainable LoRAs [24]. The joint connection module facilitates the aligned joint generation. We propose adaptive scheduling weights and an unbalanced timestep sampling strategy for effective training using separate timesteps. MM-DiT and P-DiT represent the multi-modal diffusion transformer [16] and parallel diffusion transformer [15].
@@ 2 Preliminaries
[…]
@@
However, it is intractable to obtain the true vector field $v_{t}$. To address this intractability, Lipman et al[35] proposed Conditional Flow Matching (CFM), which introduces an extra condition by sampling the accessible data sample $x_{1}$ from unknown data distribution $q(x_{1})$. By letting the true vector field conditioned on $x_{1}$, that is $v_{t}(x|x_{1})$, a tractable objective is obtained, as follows:
@@ 3.1 Joint Diffusion Transformer (JointDiT)
Figure 3 shows JointDiT architecture. JointDiT is built on Flux [5], an advanced diffusion transformer model that consists of multi-modal diffusion transformer (MM-DiT) and parallel diffusion transformer (P-DiT) blocks [16, 15]. To harness its strong image prior and the benefits of transformer architectures in dense prediction tasks [43], we extend it to joint image and depth distribution modeling by introducing a parallel depth branch alongside the pre-trained RGB branch. Thereafter, we add LoRAs [24] to the MM-DiT and P-DiT blocks to process the depth domain, which has a different data distribution from images. Additionally, joint connection modules are introduced in each DiT block to model joint distribution by interchanging features between the RGB and depth branches. We train the LoRAs and joint connection modules while keeping the pre-trained backbone model frozen.
In the joint connection modules (see Fig 3-b), feature exchange for joint distribution modeling occurs within the attention mechanism of each DiT block. We adopt the joint cross-attention module from the prior work [31]. This module facilitates joint distribution training by exchanging queries between the RGB and depth branches through attention mechanisms. The motivation is that self-attention plays a key role in the form and structure of images [54].
@@
Figure 4: 3D lifting results of LDM3D [53], JointNet [60], and our method. For each method, we lift the jointly generated image by using the jointly generated depth map. Our method generates highly plausible image-aligned 3D structures, surpassing previous joint generation methods in achieving superior consistency with real 3D space.
@@
[…]
@@ 4 Experiments
AbsRel $\downarrow$ $\delta 1$uparrow AbsRel $\downarrow$ $\delta 1$uparrow AbsRel $\downarrow$ $\delta 1$uparrow AbsRel $\downarrow$ $\delta 1$uparrow AbsRel $\downarrow$ $\delta 1$uparrow
@@
Depth-Anything-V2 [59] 4.4 97.9 7.5 94.8 6.5 95.4 13.2 86.2
@@
Table 1: Depth estimation results. We compare our method with generative joint generation methods, as well as with discriminative and generative depth estimation methods on standard depth evaluation datasets. Our approach significantly outperforms JointNet and UniCon. Additionally, it achieves comparable performance to generative depth estimation methods, except on the ETH3D dataset. Bold indicates the best performance among the generative methods in this table.
Implementation details
To collect the training dataset, we randomly sample frames from a real-world internal video dataset, which allows us to acquire real-world images with a larger field of view easily. The sampled frames are resized while maintaining their aspect ratio, then center-cropped, to produce $512\times 512$ images. The depth maps and text prompts are generated by Depth-Anything-v2 [59] and LLaVa [36], respectively. We train our model on the collected dataset, which consists of 50k pairs, for 75k iterations with a batch size of 4 and a learning rate of 1e-5. We consistently use the LoRA rank of 64 and apply text drop with a probability of 10% [22]. The training is conducted on a single NVIDIA H100 GPU for 3.5 days. More details can be found in the supplementary material.
4.1 Joint Generation
We demonstrate JointDiT’s joint generation results by showing the jointly generated images, depth maps, and their 3D lifting results, which provide the simplest way to access joint generation capability highlighting the joint distribution modeling necessity. To obtain the 3D lifting results, we apply an inverse projection to the generated image by utilizing the generated depth map. We compare our method with LDM3D [53] and JointNet [60], as they provide raw depth maps that facilitate 3D lifting visualization.
Figure 5: Depth estimation results of joint generation models. We visualize the depth estimation results of JointNet, Unicon, and our method on the NYUv2 and ScanNet dataset [14, 49]. Our approach captures thin and fine-grained details with only a timestep adjustment, i.e., $t_{x}=1$ and $t_{y}=0$. In contrast, JointNet requires additional fine-tuning for depth estimation.
Figure 4 demonstrates the results. Compared to LDM3D and JointNet, our JointDiT shows high-fidelity images, fine-detailed depth maps, and geometrically accurate 3D lifting results. In contrast, the 3D lifting results of JointNet and LDM3D are geometrically inaccurate. We assume that this significant gap in geometric accuracy is caused by the differences in the image prior and the architecture between the baseline models, i.e., stable diffusion [45] and Flux [5]. The Flux model, which is built on the diffusion transformer architecture, demonstrates extraordinary image generation quality over stable diffusion that adopts the UNet-based architecture. In addition, the transformer architecture has been shown effective in depth estimation by several studies [43, 33, 1] since it has the global receptive field different from the fully-convolutional networks. We assume this property is also effective in joint generation.
@@ 4.2 Depth Estimation
[…]
@@
Adaptive Unbalanced ImageNet 6K Pexel 6K MSCOCO 30K
@@
weights Rank1$\uparrow$ Rank2 Rank3 Rank4$\downarrow$
@@ 4.3 Depth-Conditioned Image Generation
We validate the depth-conditioned image generation quality, another joint generation with different time steps, $t_{x}=0$ and $t_{y}=1$. We compare our method with Readout-Guidance [38], ControlNet [61], and UniCon [31] using the evaluation setup proposed by UniCon. They selected methods supporting depth-conditioned image generation and trained these methods on the same training dataset, which includes 16k images of PascalVOC [17], depth maps from Depth-Anything-V2 [59], and text prompts extracted using BLIP2 [30]. For the evaluation, using the selected 6k images from the OpenImages dataset [29], they estimated depth maps using an off-the-shelf model and generated images conditioned on these depth maps and text prompts from BLIP2. For a fair comparison, we train our model on the PascalVOC dataset, following the same data processing pipeline. The evaluation is based on the FID score between the generated and original images, as well as the consistency of depth estimation results, e.g., AbsRel. Table 2 shows the depth-conditioned image generation results on the OpenImages 6K dataset. Compared to other methods, our method shows a lower FID score and AbsRel. The lower AbsRel indicates that the generated images accurately preserve the original image’s geometry.
@@ 4.4 Ablation Studies
We investigate the effectiveness of the proposed techniques for solid joint distribution modeling, that is, adaptive scheduling weights dependent on the noise levels of each modality and unbalanced timestep sampling. For comparison, we train four models on our dataset for 75k iterations, varying the use of adaptive scheduling weights and unbalanced timestep sampling by either applying or omitting. When not using unbalanced timestep sampling, we respectively sample $t_{x}$ and $t_{y}$ from timestep distribution $f(t)$, which is the distribution that our base training code suggests. First, we investigate the effect of adaptive scheduling weights and unbalanced timestep sampling on joint generation. Following the evaluation protocol of JointNet [60], we assess the quality of generated images given text prompts using 30k sampled images from the MSCOCO dataset [34]. Additionally, we construct two evaluation datasets by selecting 6k samples from ImageNet [46] and Pexel [40], respectively. For the ImageNet and Pexel datasets, where text prompts are not provided, we extract text prompts using BLIP2 [30]. We use the Inception Score (IS) [47], Fréchet Inception Distance (FID) [21], and CLIP similarity [41] as evaluation metrics. Table 3 shows that the usage of adaptive scheduling weights significantly improves all metrics across all evaluation datasets. When unbalanced timestep sampling is applied together, IS and CLIP scores tend to improve. These results suggest that considering the relative noise level is important for effectively connecting different modality generation branches during separate timestep training.
We also evaluate each model on depth estimation and depth-conditioned image generation to assess the joint distribution modeling performance at the most extreme timesteps, specifically at $t_{x}=0$ and $t_{y}=1$, and vice versa. We use the NYUv2, ScanNet, and OpenImages 6K datasets to evaluate the performance of depth estimation and depth-conditioned image generation. Table 4 reports the ablation results. In depth estimation, applying either adaptive scheduling weights or unbalanced timestep sampling improves performance. The best results are achieved when both are used together.
@@
In depth-conditioned image generation, we obtain the depth map employing Depth-Anything-V2 from the OpenImages 6K dataset, followed by depth-conditioned image generation using the estimated depth maps and text prompts obtained by BLIP2. First, we compute the FID score between the original images of OpenImage 6k and depth-conditioned generated images. As shown in Tab 4, applying only adaptive scheduling weights results in lower performance in terms of FID. However, when combined with unbalanced timestep sampling, the performance becomes comparable to other configurations. To evaluate the generated image quality beyond the distribution of the OpenImages 6K dataset, we measure the ranking measured by ImageReward [58]. ImageReward was trained by human annotators to learn human preferences in text-to-image generation. We measure the human preference ranking of images generated by each model using the same text prompt and present the corresponding percentage for rankings 1 through 4. When adaptive scheduling weights and unbalanced timestep sampling are applied together, ranking 1 has the highest proportion, while the last ranking has the lowest. These results demonstrate that the two techniques effectively model the joint distribution at extreme timesteps.

### v3

@@ Abstract
We present JointDiT, a diffusion transformer that models the joint distribution of RGB and depth. By leveraging the architectural benefit and outstanding image prior of the state-of-the-art diffusion transformer, JointDiT not only generates high-fidelity images but also produces geometrically plausible and accurate depth maps. This solid joint distribution modeling is achieved through two simple yet effective techniques that we propose, namely, adaptive scheduling weights, which depend on the noise levels of each modality, and the unbalanced timestep sampling strategy. With these techniques, we train our model across all noise levels for each modality, enabling JointDiT to naturally handle various combinatorial generation tasks, including joint generation, depth estimation, and depth-conditioned image generation by simply controlling the timesteps of each branch. JointDiT demonstrates outstanding joint generation performance. Furthermore, it achieves comparable results in depth estimation and depth-conditioned image generation, suggesting that joint distribution modeling can serve as a viable alternative to conditional generation. The project page is available at https://byungki-k.github.io/JointDiT/
Figure 1: We present JointDiT, a diffusion transformer modeling the RGB-Depth joint distribution. By leveraging the image prior of a state-of-the-art diffusion transformer [7], JointDiT generates high-fidelity images and geometrically plausible and accurate depth maps.
[…]
Figure 2: 3D lifting results of JointDiT and depth diffusion models, i.e., Marigold [32] and GeoWizard [21]. JointDiT also shows more plausible 3D point clouds than depth estimation models in challenging illustration domains, likely due to the complementary behavior of the RGB and depth branches across generative processes, i.e., the RGB branch focuses on texture, and the depth branch on structure.
@@ 1 Introduction
In the era of generative AI, diffusion models have made remarkable advancements in synthesizing images [26, 52]. The outstanding capability of text-to-image diffusion models has been found to be useful not only for image generation but also for solving important inverse problems [13, 14, 12, 28], image inpainting [15, 44], image editing [31, 16], and even further cross-modal conditional generation, such as depth-conditioned image generation [72, 6, 45], and image-conditioned depth estimation [32, 21, 22]. These works have shown that using the image prior of diffusion models is effective for modeling conditional distribution.
Recently, in image and depth modalities, joint distribution modeling [60, 71, 35] has shown that it not only enables joint generation but also shows potential as a viable alternative to existing depth estimation methods and depth-conditional image generation methods within a single unified framework by treating them as special cases of joint distribution modeling. It demonstrates that joint modeling can be easily generalized for various tasks including controllable and conditional generation and estimation. Despite this versatility, the realism of the generation is limited.
@@
To achieve this, we model the joint distribution by harnessing the strong image prior of a state-of-the-art diffusion transformer [7] and building a parallel depth branch through joint connection modules. By training on separate noise levels for each modality, JointDiT flexibly facilitates combinatorial tasks of image and depth by simply controlling the timestep of each branch. To enable separate noise level training, we propose two simple yet effective techniques, i.e., adaptive scheduling weights and unbalanced timestep sampling strategy, designed for multi-modal diffusion training with separate noise levels. JointDiT achieves significantly superior joint generation results compared to previous joint generation methods [60, 71, 35] while demonstrating comparable performance in conditional generation tasks, such as depth estimation and depth-conditioned image generation. JointDiT also enables plausible depth generation even for challenging domains such as cartoon images and pixel art illustrations, where depth estimation methods [21, 32] often struggle, as shown in Fig. 2. We further observe that the RGB and depth branches adopt complementary behavior in the joint generation process, with the depth branch capturing structural information and the RGB branch focusing on complementary aspects related to texture and appearance, which may underlie the plausible results observed in challenging domains.
@@
  • We present JointDiT, a model for solid joint distribution modeling between image and depth modalities across all noise levels by leveraging the image prior of diffusion transformers. It supports combinatorial tasks, such as joint generation, depth estimation, and depth-conditioned image generation via simple timestep control.
  • We propose adaptive scheduling weights and unbalanced timestep sampling strategy for separate noise level training in multi-modality, which significantly improves performance on combinatorial tasks. Through these techniques, we demonstrate that joint distribution modeling is a viable alternative to conditional generation.
Figure 3: Overall pipeline of JointDiT. Building on Flux [7], we introduce a parallel depth branch with trainable LoRAs [27]. The joint connection module enables the aligned joint generation. We propose adaptive scheduling weights and an unbalanced timestep sampling strategy for effective training using separate timesteps. MM-DiT and P-DiT denote the multi-modal [19] and parallel [18] diffusion transformers, respectively. The two timestep distributions, $f(t)$ and $g(t)$, are described in the supplementary material.
@@ 2 Preliminaries
Flow matching is generative modeling that learns a time-dependent vector field, which transports one probability distribution to another. It is closely related to Continuous Normalizing Flows (CNFs) [11], which model such transformations via differential equations. We adopt the notation from Lipman et al[39] to describe the flow matching formulation and objective. Given the data points $x\in\mathbb{R}^{d}$ and probability density path $p:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{>0}$ that is a time-dependent probability density function satisfying $\int p_{t}(x)dx=1$, the time-dependent vector field $v:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ combines with a flow $\phi:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, leading to the ordinary differential equation (ODE):
@@
However, obtaining the true vector field $v_{t}$ is intractable. To address this, Lipman et al[39] proposed Conditional Flow Matching (CFM), which introduces a condition by sampling the accessible data sample $x_{1}$ from the unknown data distribution $q(x_{1})$. By conditioning the true vector field on $x_{1}$, that is $v_{t}(x|x_{1})$, a tractable objective is obtained, as follows:
@@ 3.1 Joint Diffusion Transformer (JointDiT)
Figure 3 shows JointDiT architecture. JointDiT is built on Flux [7], an advanced diffusion transformer model that consists of multi-modal diffusion transformer (MM-DiT) and parallel diffusion transformer (P-DiT) blocks [19, 18]. To harness its strong image prior and the benefits of transformer architectures in dense prediction tasks [50], we extend it to joint image and depth distribution modeling by introducing a parallel depth branch alongside the pre-trained RGB branch. Thereafter, we add LoRAs [27] to the MM-DiT and P-DiT blocks to process the additional depth domain. Additionally, joint connection modules are introduced in each DiT block to model joint distribution by interchanging features between the RGB and depth branches. We train the LoRAs and joint connection modules while keeping the pre-trained backbone model frozen.
In the joint connection modules (see Fig. 3-b), feature exchange for joint distribution modeling occurs within the attention mechanism of each DiT block. We adopt the joint cross-attention module from the prior work [35]. This module facilitates joint distribution training by exchanging queries between the RGB and depth branches through attention mechanisms. The motivation is that self-attention plays a key role in the form and structure of images [63].
@@
Figure 4: 3D lifting results of LDM3D [60], JointNet [71], and our JointDiT. Our JointDiT generates highly plausible image-aligned 3D structures, surpassing previous joint generation methods in achieving superior consistency with real 3D space.
@@
We also introduce the unbalanced timestep sampling strategy to enforce joint distribution modeling at any separate timesteps (See Fig. 3-c). Prior studies [19, 73] have investigated the impact of timestep sampling strategies on diffusion performance during training and have proposed various timestep sampling methods beyond uniform distribution. Similarly, our base training code also employs a weighted timestep distribution for training [^2] https://github.com/kohya-ss/sd-scripts/tree/sd3 ($f(t)$ in Fig 3-c). However, with this timestep distribution, the joint distribution of $t_{x}$ and $t_{y}$ is likely insufficient to fully cover both joint generation and conditional generation tasks, as shown by Hang et al[23], where insufficient timestep sampling negatively affected diffusion performance. The unbalanced timestep sampling strategy samples $t_{x}$ and $t_{y}$ independently from two unbalanced timestep distributions, $f(t)$ and $g(t)$, with half probability during training. For the remaining half, the same timesteps sampled from $f(t)$ are assigned to $t_{x}$ and $t_{y}$. We experimentally validate that these two simple techniques are effective for building a solid joint distribution across all noise levels of images and depth maps, enhancing performance in joint generation, depth-conditioned image generation, and depth estimation.
@@ 4 Experiments
We evaluate the performance of JointDiT across joint generation, depth estimation, and depth-conditioned image generation. We also analyze the behavior of the RGB and depth branches, as well as the effectiveness of adaptive scheduling weights and unbalanced timestep sampling. Experimental details can be found in the supplementary material.
Implementation details
To collect the training dataset, we randomly sample frames from a real-world internal video dataset, which allows us to acquire real-world images with a larger field of view easily. The sampled frames are resized while maintaining their aspect ratio, then center-cropped, to produce $512\times 512$ images. The depth maps and text prompts are generated by Depth-Anything-v2 [68] and LLaVa [40], respectively. We train our model on the collected dataset, which consists of 50k pairs, for 75k iterations with a batch size of 4 and a learning rate of 1e-5. We use the LoRA rank of 64 in DiT blocks and apply text drop with a probability of 10% [25]. The training is conducted on a single NVIDIA H100 GPU for 3.5 days.
4.1 Joint Generation
We demonstrate JointDiT’s joint generation capability through visualizations of generated images, depth maps, and their 3D lifting results, which provide intuitive evidence of joint RGB–Depth modeling and underscore the necessity of joint distribution modeling. To obtain the 3D lifting results, we apply an inverse projection to the generated image using the generated depth map. We compare our method with LDM3D [60] and JointNet [71], as they provide raw depth maps that facilitate 3D lifting visualization.
Figure 4 demonstrates the results. Compared to LDM3D and JointNet, our JointDiT shows high-fidelity images, fine-detailed depth maps, and geometrically accurate 3D lifting results. In contrast, the 3D lifting results of JointNet and LDM3D are geometrically inaccurate. We assume that this significant gap in geometric accuracy is caused by the differences in the image prior and the architecture between the baseline models, i.e., stable diffusion [52] and Flux [7]. The Flux model, which is built on the diffusion transformer architecture, demonstrates superior image generation quality over stable diffusion that adopts the UNet architecture. In addition, the transformer architecture has been shown effective in depth estimation by several studies [50, 37, 1] since it has the global receptive field different from the fully-convolutional networks.
@@ 4.1 Joint Generation
AbsRel $\downarrow$ $\delta 1$ $\uparrow$ AbsRel $\downarrow$ $\delta 1$ $\uparrow$ AbsRel $\downarrow$ $\delta 1$ $\uparrow$ AbsRel $\downarrow$ $\delta 1$ $\uparrow$ AbsRel $\downarrow$ $\delta 1$ $\uparrow$
@@
Depth-Anything-V2 [68] 4.4 97.9 4.1 97.9 7.5 94.8 6.5 95.4 13.2 86.2
4M-21 [3] 11.8 88.7 10.6 89.1 15.6 78.9 32.1 75.1 8.4 93.8
MultiMAE [2] 9.2 91.6 8.7 92.3 16.9 75.1 35.2 71.9 10.6 89.9
Unified-IO [42] 6.8 95.9 7.5 95.0 28.1 52.0 36.4 70.0 13.9 83.9
Unified-IO 2 [43] 12.5 85.8 14.5 81.6 48.9 31.7 43.4 62.0 20.5 72.1
@@
Table 1: Depth estimation results. We compare ours with generative joint generation methods, as well as with discriminative and generative depth estimation methods. Ours outperforms JointNet and UniCon. Additionally, it achieves comparable performance to generative depth estimation methods, except on ETH3D. Bold indicates the best performance among the generative methods in this table.
@@ 4.2 Depth Estimation
We assess the depth estimation capability of JointDiT with different time steps, $t_{x}=1$ and $t_{y}=0$. We compare our method with joint generation methods that support depth estimation, e.g., JointNet [71] and UniCon  [35]. We also compare with discriminative depth estimation methods, including depth-specialized models [49, 50, 68] and multitask learning-based methods [3, 2, 42, 43], as well as generative depth estimation methods that utilize a diffusion model [32, 21]. Following the evaluation convention of prior work, we compare each method on the NYUv2 [56], ScanNet [17], KITTI [5], DIODE [64], and ETH3D [55] datasets. The evaluation metrics are Absolute Mean Relative Error (AbsRel) and $\delta_{1}$. Table 1 summarizes the results. Compared to joint generation methods, our model achieves superior performance across all evaluation datasets. Figure 5 shows the depth estimation results of joint generative methods on the ScanNet dataset. Compared to JointNet and UniCon, our method captures sharp edges and fine details.
@@
Figure 5: Depth estimation results of joint generation models. We visualize the depth estimation results of JointNet, UniCon, and our method on the NYUv2 and ScanNet dataset. Our approach captures thin and fine-grained details with only a timestep adjustment, i.e., $t_{x}=1$ and $t_{y}=0$. In contrast, JointNet requires additional fine-tuning for depth estimation.
@@
Adaptive Unbalanced ImageNet 6K Pexels 6K MSCOCO 30K
@@
[…]
@@ 4.3 Depth-Conditioned Image Generation
We validate the depth-conditioned image generation quality, another joint generation with different time steps, $t_{x}=0$ and $t_{y}=1$. Following the evaluation protocol of UniCon [35], We compare our method with Readout-Guidance [45], ControlNet [72], and UniCon. All methods are trained on the same dataset, i.e., 16k samples from PascalVOC [20], and evaluated on 6k samples from OpenImages [33]. The evaluation is based on the FID score between the generated and original images, as well as the consistency of depth estimation results, e.g., AbsRel. Table 2 shows the depth-conditioned image generation results on the OpenImages 6K dataset. Compared to other methods, our method shows a lower FID score and AbsRel. The lower AbsRel indicates that the generated images accurately preserve the original image’s geometry.
Figure 6: Feature visualization of RGB-Depth branches in $M$-DiT blocks. We observe that, at timestep $t=0.48$, the depth branch tends to focus on scene geometry, while the RGB branch captures semantic patterns related to texture and appearance.
@@ 4.4 Ablation Studies
Joint RGB-Depth feature visualization
As shown in Fig. 2, joint RGB-Depth modeling enables plausible depth generation from text prompts, even for stylized domains such as cartoon-style and pixel art illustrations. Furthermore, it tends to yield more structured and plausible 3D lifting results, not only for stylized domains but also for realistic image domains, compared to depth prediction from RGB images alone, as can be found in the supplementary material. The way the RGB and depth branches focus on textures and structure, respectively, may hint at how such capability achieves. Figure 6 shows the joint RGB-Depth features of MM-DiT blocks, visualized following the method of Tumanyan et al[63]. The depth branch focuses on capturing underlying structural information, while the RGB branch attends to complementary aspects related to texture and appearance. This behavior may suggest that the depth branch, by focusing on geometric properties throughout the joint generation process, enables plausible depth generation even in scenes where depth estimation is challenging.
Adaptive scheduling weights and unbalanced timestep sampling
@@
We analyze the effectiveness of adaptive scheduling weights and unbalanced timestep sampling in achieving solid joint distribution modeling. For comparison, we train four models on our dataset for 75k iterations, varying the use of adaptive scheduling weights and unbalanced timestep sampling by either applying or omitting. When not using unbalanced timestep sampling, we respectively sample $t_{x}$ and $t_{y}$ from timestep distribution $f(t)$, which is the distribution that our base training code suggests. We first investigate the effect in joint generation by assessing the quality of images generated from text prompts on three datasets: 30k samples from MS-COCO [38], 6k from ImageNet [53], and 6k from Pexels [47]. As evaluation metrics, we use the Inception Score (IS) [54], Fréchet Inception Distance (FID) [24], and CLIP similarity [48] as evaluation metrics. Table 3 shows that the usage of adaptive scheduling weights significantly improves all metrics across all evaluation datasets. When unbalanced timestep sampling is applied together, IS and CLIP scores tend to improve. These results suggest that considering the relative noise level is important for effectively connecting different modality generation branches in separate timestep training.
@@
We also evaluate each model on depth estimation and depth-conditioned image generation to assess the joint distribution modeling performance at the most extreme timesteps, specifically at $t_{x}=0$ and $t_{y}=1$, and vice versa. We use the NYUv2, ScanNet, and OpenImages 6K datasets to evaluate the performance of depth estimation and depth-conditioned image generation. Table 4 reports the results. In depth estimation, applying either adaptive scheduling weights or unbalanced timestep sampling improves performance. The best results are achieved when both are used together. For the depth-conditioned image generation, we follow UniCon [35] for the evaluation setting, using 6K samples from the OpenImages dataset. We report the FID between original and depth-conditioned images, and also the percentage of samples ranked 1st to 4th by ImageReward [67], a human preference-trained reward model for text-to-image generation. As shown in Tab 4, applying only adaptive scheduling weights results in lower performance in terms of FID. However, when combined with unbalanced timestep sampling, the performance becomes comparable to other configurations. Notably, when adaptive scheduling weights and unbalanced timestep sampling are applied together, ImageReward ranking 1 has the highest proportion, while the last ranking has the lowest. These results demonstrate that the two techniques effectively model the joint distribution at extreme timesteps.